Sandhill, Inc., has outstanding bonds that will mature in six years and pay an 8 percent coupon semiannually. If you paid $1,084.29 today and your required rate of return was 5.6 percent. (Round intermediate calculations to 5 decimal places, e.g. 1.25145 and final answer to 2 decimal places, e.g. 15.25.)

How much should you have paid for the bond?

Worth of the bond $______

Did you pay the right price for the bond? Good, Fair, or Bad
Answer #1

Bond Valuation: The value of bond is the present value of the expected cashflows from the bond,discounted at Yield to Maturity(YTM).

Year Cash flow PVAF/PVF@2.8% Present Value (Cashflow*PVAF/PVF)
1-12 40 10.0739* 402.96
12 1000 0.71793** 717.93

Current Market Price of Bonds = 402.96+717.93

= 1120.89

*PVAF = (1-(1+r)^-n)/r

**PVF = 1 / (1+r)^n

Note : Since the bond makes semiannual interest payments, total no. of period is 12 (6*2), cashflow per period is 40(1000*8%/2) and cashflows are discounted at 2.8% (5.6/2).

Since the fair market price (1120.89) > cost of acquisition (1084.29) ,  pay the Good price for the bond